Mrs Ada is three times as old as her daughter.six years ago,the ratio of their ages were 5:1.find her daughters age
Answer by stanbon(75887)

Mrs Ada is three times as old as her daughter.six years ago,the ratio of their ages were 5:1.find her daughters age
------
Equations:
a = 3d
(a-6)/(d-6) = 5
------------------------
Substitute for "a" and solve for d.
--------
(3d-6) = 5(d-6)
3d-6 = 5d-30
2d = 24
d = 12 (daughter's age now)
----
a = 3d = 36 (Mother's age now)
-----------
Cheers,
Stan H.
-------------